Steve's Auto Shop began the year with an inventory of $12,050. Purchases during the year totaled $70,600. The inventory at the end of the year was $13,530.

a. What was the cost of goods sold?

$

b. What was the average inventory?

$

c. What was the turnover? Round your answer to two decimal places.
Answer #1

Answer

--Requirement a

A Beginning Inventory $12,050
B Purchases $70,600
C Ending Inventory $13,530
D = A+B - C Cost of Goods Sold $69,120

--Requriement b

A Beginning Inventory $12,050
B Ending Inventory $13,530
C = A+B Total $25,580
D = C/2 Average Inventory $12,790

--Requirement c

A Cost of Goods Sold $69,120
B Average Inventory $12,790
C = A/B Turnover 5.40